OpenStudy (anonymous):

8/7 divided by 3/7 i need steps to solve

OpenStudy (anonymous):

\(\dfrac{8}{7} \div \dfrac{3}{7}\) Try thinking of it this way.... How many \(\dfrac{3}{7}\) portions can you get out of \(\dfrac{8}{7}\)? The steps will be: 1. Same - keep the first fraction the same. 2. Change - change division to multiplication. 3. Flip - flip the second fraction.

OpenStudy (anonymous):

24/49 final answer or can i simplify anymore i think not

OpenStudy (anonymous):

I think you forgot to flip...

OpenStudy (anonymous):

56/21

OpenStudy (anonymous):

That can be simplified. But to keep the numbers smaller, I like to cancel before I multiply.

OpenStudy (anonymous):

\(\dfrac{8}{\cancel{7}} \times \dfrac{\cancel{7}}{3} = \dfrac{8}{3}\)

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

it does make sense that was pretty basic thank u

OpenStudy (anonymous):

Your welcome.